OpenStudy (anonymous):

a catapult lunches a boulder with an upword velocity of 112ft/sec. the height of the boulder 'h' in feet after 't' second is given by the function h = -16t^2+ 112t +30. how long does it take the boulder to reach its maximum height?what is the boulder maximum height? round to the nearest hundredth if necessary.

OpenStudy (anonymous):

any one can you help to solve this?

OpenStudy (whpalmer4):

Okay, do you recognize what sort of equation h = -16t^2 + 112t + 30 is?

OpenStudy (anonymous):

it is qudratic one

OpenStudy (whpalmer4):

okay, what sort of curve does it make?

OpenStudy (anonymous):

i think it should be upword

OpenStudy (whpalmer4):

parabola, hyperbola, circle, ellipse?

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

don't think of the curve as the trajectory of the boulder, this is a plot of height vs. time...

OpenStudy (whpalmer4):

it's a parabola.

OpenStudy (anonymous):

thank you

OpenStudy (whpalmer4):

does this parabola open upward (looks like a bowl) or does it open downward (upside down bowl)?

OpenStudy (whpalmer4):

hint: the problem asks for maximum height :-)

OpenStudy (anonymous):

maximum height intial velocity is 0

OpenStudy (whpalmer4):

where did you get the idea that the initial velocity is 0?

OpenStudy (whpalmer4):

before launch, it is 0, but at the time of launch, it is 112 ft/s, upward

OpenStudy (whpalmer4):

that's represented by the 112t term in the equation

OpenStudy (whpalmer4):

the -16t^2 term represents the effects of gravity accelerating the boulder toward the ground. the +30 term represents the height at the time of launch

OpenStudy (anonymous):

physics

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

yeah, but you don't actually need to know any physics to solve this problem. just a little algebra and analytic geometry

OpenStudy (whpalmer4):

because the problem asks for the maximum height, this must be a downward opening parabola. if the parabola opens upward, there is no maximum height, right?

OpenStudy (anonymous):

i thinks when parabola is downword it is maximum ya?

OpenStudy (whpalmer4):

yes, the maximum will be at the vertex of the parabola. do you know how to find the vertex from that equation?

OpenStudy (anonymous):

please show me

OpenStudy (whpalmer4):

okay, there are a number of ways to do it. what class are you taking?

OpenStudy (anonymous):

college algebra 1

OpenStudy (anonymous):

ple show me the easyest way

OpenStudy (whpalmer4):

okay, then we won't use calculus :-) so to find the vertex of a parabola we can rearrange the formula into "vertex form" which is y = a(x-h)^2 + k with the formula in vertex form, we can read off the coordinates of the vertex directly: it is simply (h, k)

OpenStudy (whpalmer4):

here we have height = -16t^2 + 112t + 30 we need to rearrange that as height = -a(t - h)^2 + k any idea how to do that? have you heard of completing the square?

OpenStudy (anonymous):

ya but need help

OpenStudy (anonymous):

(b/2)^2

OpenStudy (anonymous):

is the formula?

OpenStudy (whpalmer4):

first thing to do is factor out the -16 height = -16(t^2 - (112/16)t - 30/16) then b = -112/16 and b/2 = -56/16 = -7/2 and (b/2)^2 = 49/4 agreed?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thank you

OpenStudy (whpalmer4):

that gives us height = -16(t-7/2)^2 +16*(49/4)+30 or height = -16(t-7/2)^2 + 226

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

h = 7/2 k = 226 so the vertex is at (7/2, 226) or (3.5, 226)

OpenStudy (anonymous):

thank you very much

OpenStudy (whpalmer4):

a different approach is to remember that if you have a parabola in standard form y = ax^2 + bx + c (which we do, without any rearrangement) then the x-coordinate of the vertex is at x = -b/(2a) and the y-coordinate of the vertex is just the function evaluated at that value of x so we will do it that way as well, as a check on our first solution: height = -16t^2 +112t + 30 a = -16 b=112 c = 30 vertex x-coordinate = - b/(2a) = -112/(2*-16) = -112/-32 = 56/16 = 7/2 = 3.5 vertex y-coordinate = -16(3.5)^2+112(3.5) + 30 = -16(12.25) + 112(3.5) + 30 = 226

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

Here's a graph:

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i got it now thank you very much for your great help

OpenStudy (whpalmer4):

Glad I could help! Thanks for the medal, you just promoted me a rank :-)

OpenStudy (anonymous):

you help me a lot thanks for your idea have a good night

OpenStudy (whpalmer4):

you too!